OpenStudy (anonymous):

(7,1) perpendicular to y=-x+3

OpenStudy (gorv):

line perpendicular to y=-x+3 slope1=-1 for perpendicular line slope1*slope2=-1 slope2=-1/slope1=-1/-1=-

OpenStudy (gorv):

line having slope 1 y=x+c and it pass through (7,1) 1=7+c c=-6 y=x-6